OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=8, b=4, c=?

OpenStudy (anonymous):

so do you know the formula?

OpenStudy (anonymous):

a^2 + B^2=c^2

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

oh

OpenStudy (purplemexican):

8^2+4^2=c^2 youll want to \[\sqrt{c}\]

OpenStudy (anonymous):

do whaa

OpenStudy (anonymous):

so just plug a and b in and then you take the square root of c

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

so 80?

pooja195 (pooja195):

This is called the Pythagorean theorem \[\huge~\rm~a^2+b^2=c^2\]  \[\huge~\rm~8^2+4^2=c^2\] \[\huge~\rm~64+16=c\]

pooja195 (pooja195):

Once you find 64+16 just square root it

OpenStudy (anonymous):

yes then you take the square root of that

OpenStudy (anonymous):

i dont know how to square root haha that is why i need help

pooja195 (pooja195):

whats 64+16?

OpenStudy (purplemexican):

\[8^2+4^2=c^2\ or \sqrt{80}\]

pooja195 (pooja195):

right you might need to use a calc for that when you do you get a decimal \[\huge~\rm~8.944272 \]

pooja195 (pooja195):

From there just take the first 3 numbers

pooja195 (pooja195):

so about 8.94

OpenStudy (anonymous):

or you could simplify 80.

OpenStudy (anonymous):

so 9?

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 so about 8.94 \(\color{#0cbb34}{\text{End of Quote}}\) this

OpenStudy (anonymous):

guys this is a multiple choice and the answers aern't numbers like that

OpenStudy (anonymous):

\[\sqrt{16+5}=4\sqrt{5}\]

OpenStudy (anonymous):

that is the answer

pooja195 (pooja195):

@kylefoster58 it would help if u put the answer choices

OpenStudy (anonymous):

sorry haha

OpenStudy (lfreeman1285):

Here you use the Pythagorean theorem which is a^2 + b^2 = c^2 a and b are the legs of the triangle while c is the hypotenuse. All you have to do is fill in the variables and solve for c. Therefore, (8)^2 + (4)^2 = c^2 64 + 16 = 80 But because 80 is c^2, you have to square root it. \[\sqrt{80}\] = \[4\sqrt{5}\] and that's your final answer.

pooja195 (pooja195):

^thats right

OpenStudy (anonymous):

its D.

OpenStudy (anonymous):

4 square root of 5 right

OpenStudy (anonymous):

thank yall again!!